Professor Wendy Smith has been offered the following​ opportunity: A law firm would like to retain her for an upfront payment of $49,000. In​ return, for the next year the firm would have access to eight hours of her time every month. As an alternative payment​ arrangement, the firm would pay Professor​ Smith's hourly rate for the eight hours each month. Smith's rate is $535 per hour and her opportunity cost of capital is 15% per year. What does the IRR rule advise regarding the payment​ arrangement? (Hint: Find the monthly rate that will yield an effective annual rate of 15%.) What about the NPV​ rule?
Answer #1
Assume monthly rate =i
(1+i)^12=(1+0.15)
1+i=1.15^(1/12)= 1.011715
i=monthly interest rate 0.011715
Present Value of cash flow:
(cash flow)/((1+i)^N)
i=discount rate =0.011715
N=Month of Cash Flow
CASH FLOW :
If the Second option (monthly payment) is chosen
Opportunity cost in month 0 ($49,000)
Cash Inflow in every month =8*535 $4,280
N CF PV=CF/(1.011715^N)
Month Cash Flow Present Value
0 ($49,000) ($49,000)
1 $4,280 $4,230.44
2 $4,280 $4,181.45
3 $4,280 $4,133.04
4 $4,280 $4,085.18
5 $4,280 $4,037.87
6 $4,280 $3,991.12
7 $4,280 $3,944.90
8 $4,280 $3,899.22
9 $4,280 $3,854.07
10 $4,280 $3,809.45
11 $4,280 $3,765.34
12 $4,280 $3,721.74
SUM ($1,346)
NPV =SUM of PV= ($1,346)
IRR=Internal rate of return -0.44% (Using IRR function of excel over Cash Flow)
IRR rule advises to take lump sum payment
NPV rule advises to take lump sum payment
LIPDUDIU 91 PUPIL Allgrimieril Number D31 EIRR(E16:E28) B E 16 17 19 20 21 22 0 ($49,000) 1 $4,280 2 $4,280 $4,280 4 $4,280 5